State which of the following are not a function from R to R and why.

(a)

f(x) = 1/x

(b)

f(x) = 1/(1+x)

(c)

f(x) = (x)½

(d)

f(x) = ±(x2+1)½

(e)

f(x) = sin(x)

(f)

f(x) = ex


1
Expert's answer
2021-07-16T13:36:09-0400

Solution:

A relation is a function if for each value of "x", there exists a unique value of "y" or "f(x)".

For all of the followings, we have this condition satisfied:

(a) f(x) = 1/x

(b) f(x) = 1/(1+x)

(c) f(x) = (x)½

(e) f(x) = sin(x)

(f) f(x) = ex

But for (d) f(x) = ±(x2+1)½, we have two values of f(x), thus it is not a function.